OpenStudy (anonymous):

change 2/7% to a ratio

OpenStudy (cwrw238):

this equals 2 / 7/100 = ( 2 * 100) / 7

OpenStudy (dominusscholae):

(2/7)/100= (2/700)= 1/350.